PUBLICATION LIST OF JACQUES MARTINET

Last update: March 14th, 2024

Remarks. 1. Papers which account for seminar talks which would have doubled another item are note quoted in this publication list.

2. My recent publications and various texts or slides can be downloaded from my home page

http://jamartin.perso.math.cnrs.fr/ .

References

[1] J. Martinet, J.-J. Payan, Sur les extensions cubiques des rationnels et leur clôture galoisienne, J. reine angew. Math. 228 (1967), 15–37.

[2] J. Martinet, J.-J. Payan, Sur les bases d’entiers des extensions galoisiennes et non abéliennes de degré 6 des rationnels, J. reine angew. Math. 229 (1968), 29–33.

[3] J. Martinet, Sur l’arithmétique des extensions galoisiennes à groupe de Galois diédral d’ordre 2p Ann. Inst. Fourier 19,1 (1969), 1–80.

[Announced in Comptes Rendus Acad. Sc. Paris, May 6th, 1968.]

[4] J. Martinet, Anneau des entiers d’une extension galoisienne considéré comme module sur l’algèbre du groupe de Galois, Mémoire S.M.F. 25 (1971), 123–126.

[5] J. Martinet, Un contre-exemple à une conjecture d’E. Noether (d’après Swan), Séminaire Bourbaki 1969/1970, exposé ^{∘} 372 ; Springer Lecture Notes (Heidelberg) 180 (1971), 145–154.

[6] J. Martinet, Sommes de carrés, Séminaire de Théorie des Nombres de Bordeaux, 1970/1971, exposé n^{∘} 26, 9 pp.

[7] Modules sur l’algèbre du groupe quaternionien, Ann. Sc. E.N.S. 4 (1971), 399–408.

[8] J. Martinet, Sur les extensions à groupe de Galois quaternionien, Comptes Rendus Acad. Sc. Paris 274 (1972), 933–935.

[9] P. Damey, J. Martinet, Plongement d’une extension quadratique dans une extension quaternionienne, J. reine angew. Math. 262/263 (1973), 323-338.

[10] J. Martinet, Bases normales et constante de l’équation fonctionnelle des fonctions L d’Artin, Séminaire Bourbaki 1973/1974, exposé n^{∘} 450 ; Springer Lecture Notes (Heidelberg) 431 (1975),

[11] J. Martinet, Character Theory and Artin L-functions, in Algebraic Number Fields (Fröhlich ed.), Academic Press, New-York, 1977, 1–80.

[12] J. Martinet, H_{8}, same book, 525–537.

[13] J. Martinet, Tours de corps de classes et estimations de discriminants, Invent. Math. 44 (1978), 65–73.

[14] J. Martinet, Petits discriminants, Ann. Inst. Fourier 29,1 (1979), 159–170.

[15] J. Martinet, Petits discriminants des corps de nombres, Lectures Notes of the London Math. Soc. 56 (1982), 151–193.

[16] A. Leutbecher, J. Martinet, Lenstra’s constant and Euclidean number fields, Astérisque 94 (1982), 87–131. See also:

[16a] J. Martinet, Sur la constante de Lenstra des corps de nombres, Séminaire de Théorie des Nombres de Bordeaux, 1979/1980, exposé n^{∘} 17, 21 pp;

[16b] A. Leutbecher, J. Martinet, Constante de Lenstra des corps de nombres euclidiens, Séminaire de Théorie des Nombres de Bordeaux, 1981/1982, exposé n^{∘} 4, 7 pp.

[17] J. Martinet, Appendice à l’article de J-F. Mestre “Courbes elliptiques et groupes de classes d’idéaux de certains corps quadratiques”, J. reine angew. Math. 343 (1983), 32–34.

[18] J. Martinet, Préface pour la seconde édition de l’ouvrage de Hasse : Über die Klassenzahl abelscher Zahlkörper, Akademie-Verlag and Springer-Verlag, Berlin and Heidelberg, 1985, 10 pp.

[19] J. Martinet, Méthodes géométriques dans la recherche des petits discriminants, Séminaire de Théorie des Nombres de Paris 1983–84, Birkhäuser, Basel, 1985, 147–179.

[20] A.-M. Bergé, J. Martinet, Formes quadratiques et extensions en caractéristique 2, Ann. Inst. Fourier 35,2 (1985), 57–77.

[21] S-H. Kwon, J. Martinet, Sur les corps résolubles de degré premier, J. reine angew. Math. 375/376 (1987), 12–23.

[22] H. Cohen, J. Martinet, Class Groups of Number Fields: Numerical Heuristics, Math. Comp. 48 (1987), 123–137.

[23] H. Cohen, J. Martinet, Ein heuristisches Studium der Klassengruppen, Forschungsgesellschaft Joanneum Berichte, Graz (1987), 60–69.

[24] A.-M. Bergé, J. Martinet, Sur un problème de dualité lié aux sphères en géométrie des nombres, J. Number Theory 32 (1989), 14–42.

[25] A.-M. Bergé, J. Martinet, Sur les minorations géométriques des régulateurs, Séminaire de Théorie des Nombres de Paris, 1987-88, Birkhäuser, Boston, 1989, 23–50.

[26] A.-M. Bergé, J. Martinet, Notions relatives de régulateurs et de hauteurs, Acta Arith. 54 (1989), 155–170.

[27] A.-M. Bergé, J. Martinet, Minorations de hauteurs et petits régulateurs relatifs, Séminaire de Théorie des Nombres de Bordeaux, 1988–1989, exposé 11, 28 pp.

[28] H. Cohen, J. Martinet, Étude heuristique des groupes de classes, J. reine angew. Math. 404 (1990), 39–76.

[29] J. Martinet, Les discriminants quadratiques et la congruence de Stickelberger, J. Théorie des Nombres de Bordeaux 1 (1989), 197–204.

[30] A.-M. Bergé, J. Martinet, M. Olivier, The computation of sextic fields with a quadratic subfield, Math. Comp. 54 (1990), 869–884. [Known as “BéMOl”.]

[31] J. Martinet, Discriminants and permutation groups, Number Theory, R. Mollin ed., W. de Gruyter, Berlin, New-York, 1990, 359–385.

[32] F. Diaz y Diaz, J. Martinet, M. Pohst, The minimum discriminant of totally real octic fields, J. Number Theory 36 (1990), 145–159.

[33] A.-M. Bergé, J. Martinet, Réseaux extrêmes pour un groupe d’automorphismes, Astérisque 198-200 (1992), 41–66.

[34] H. Cohen, J. Martinet, Heuristics on Class Groups: Some Good Primes are not too Good, Math. Comp 63 (1994), 329–334.

[35] A.-M. Bergé, J. Martinet, F. Sigrist, Une généralisation de l’algorithme de Voronoï pour les formes quadratiques, Astérisque 209 (1992), 137–158.

[36] E. Bayer-Fluckiger, J. Martinet, Formes quadratiques liées aux algèbres semisimples, J. reine angew. Math. 451 (1994), 51–69.

[37] J. Martinet, Structures algébriques sur les réseaux, Séminaire de Théorie des Nombres de Paris, 1992-93; Number Theory, London Math. Soc. Lecture Note Series 215, S. David ed., Cambridge University Press (1995), 167–186.

[38] C. Batut, J. Martinet, Radiographie des réseaux parfaits, Experimental Math. 3 (1994), 39–49.

[39] A.-M. Bergé, J. Martinet, Sur la classification des réseaux eutactiques. J. London Math. Soc. 53 (1996), 417–432.

[40] A.-M. Bergé, J. Martinet, Densité dans des familles de réseaux. Application aux réseaux isoduaux, L’Enseignement Mathématique 41 (1995), 335–365.

[41] J. Martinet, Histoire des Journées Arithmétiques, J. Théorie des Nombres de Bordeaux 7 (1995), vii–xiii.

Updated versions are available at http://jamartin.perso.math.cnrs.fr/ .

[42] J. Martinet, Les réseaux parfaits des espaces euclidiens, Masson (now Dunod), Paris, 1996 (400 pp.). English updated edition: [57].

[43] J. Martinet, Une famille de réseaux dual-extrêmes, J. Théorie des Nombres de Bordeaux 9 (1997), 169–181.

[44] J. Martinet, Algebraic Constructions of Lattices; Isodual Lattices, Acts of the Eger congress (Hungary, 1996), Walter de Gruyter, Berlin, New York, 1997, 349–360.

[45] J. Martinet, La géométrie des nombres : une application des espaces euclidiens, Actes du congrès inter-IREM de Bayonne, juin 1996, I.R.E.M., Université Bordeaux 1, juin 1996

[46] J. Martinet, Perfect and Eutactic Lattices: Extensions of Voronoi’s Theory, in Voronoï’s impact on modern science, P. Engel, H. Syta, ’ed., Proc. Inst. Math. National Acad. Sc. Ukraine, vol. 21, Book 1 (1998), 186–198.

[47] J. Martinet, On Successive Minima of Rings of Algebraic Integers, Acts of the conference ANTS III, Portland, Oregon, 1998; Springer Lectures Notes in Computer Science 1423 (1998), 424–432.

[48] J. Martinet, Sur la classification des réseaux parfaits de dimension 5, J. Théorie des Nombres de Bordeaux (Journées Arithmétiques de Limoges de 1997)

11 (1999), 149–159. [See [57], Section 6.4.]

[49] J. Martinet, ed. Réseaux euclidiens, designs sphériques et formes modulaires, L’Enseignement Mathématique, Monographie n∘ 37, Genève, 2001, 250 pp.

[50] B. Venkov (Notes by J. Martinet), Réseaux et designs sphériques, [49], 10–86.

[After lectures by Boris Venkov, Bordeaux, 1996/1997.]

[51] J. Martinet, B. Venkov, Les réseaux fortement eutactiques (with an appendix by R. Coulangeon), [49], 112–134.

[52] J. Martinet, Sur certains designs sphériques liés à des réseaux entiers, [49], 135–146.

[53] J. Martinet, Sur l’indice d’un sous-réseau (with an appendix by C. Batut), [49], 163–211.

[54] J. Martinet, Reduction Modulo 2 and 3 of Euclidean Lattices, J. Algebra 251 (2002), 864–887.

[55] J. Martinet, The Perfection Default Module of a Euclidean Lattice; see my home page, http://math.u-bordeaux1.fr/ martinet.

[56] J. Martinet, L’arithmétique des algèbres de quaternions, notes of lectures delivered in Lausanne, May 2002; see my home page.

[57] J. Martinet, Perfect Lattices in Euclidean Spaces, Grundlehren 327, Springer-Verlag, Heidelberg, 2003, 527 pp. [English updated edition of [42].]

[58] A.-M. Bergé, J. Martinet, Symmetric Groups and Lattices, Monatshefte Math. 140 (2003), 179–195.

[59] J. Martinet, Reduction Modulo 2 and 3 of Euclidean Lattices, II (Acts of the Conference on Quadratic Forms, Talca, Chile, December 2002), Contemporary Math. 344 (2004), 265–273.

[60] J. Martinet, B. Venkov, On integral lattices having an odd minimum, Algebra i Analiz (Saint-Petersburg) 16, 3 (2004), 99–142.

[61] J. Martinet, On Parity Classes. notes of lectures delivered in Lausanne, May 2003; see my home page.

[62] A.-M. Bergé, J. Martinet, A generalization of some lattices of Coxeter, Mathematika 51 (2004), 49–61.

[63] A.-M. Bergé, J. Martinet, Sublattices of certain Coxeter lattices (A paper dedicated to Georges Gras), J. Th. Nombres Bordeaux 17 (2005), 455–465.

[64] A.-M. Bergé, J. Martinet, On weakly eutactic forms, J. London Math. Soc. 75 (2007), 187-198.

[65] A.-M. Bergé, J. Martinet, On Perfection Relations in Lattices, Contemporary Math. 493 (2009), 29–49; (Algebraic and Arithmetic Theory of Qu(a)dratic Forms, Frutilar, Chile, December 13–19, 2007); preprint: arXiv, math.NT/0611220, 8 Nov. 2006, 26 pp.

[66] J. Martinet, Bases of minimal vectors in Euclidean lattices, I , Archiv Math. 89 (2007), 404–410.

[67] J. Martinet, Bases of minimal vectors in Euclidean lattices, II , Archiv Math. 89 (2007), 541–551.

[68] J. Martinet, Anne-Marie Bergé — In Memoriam, J. Th. Nombres Bordeaux 20, fasc. 2 (2008), i–xi.

[69] W. Keller, J. Martinet, A. Schürmann, On classifying Minkowskian sublattices, Math. Comp., 81 (2012), 1063-1092; see also arXiv:0904.3110v3 [math.NT].

[70] J. Martinet, A. Schürmann, Bases of minimal vectors in Euclidean lattices, III , Inter. J. Number T. 08, 2 (2012), 551–567; see also arXiv:1105.5889v1 [math.NT].

[71] J. Martinet, On the index system of well-rounded lattices, preprint at arXiv:1202.2295v1 [math.NT] (10 Feb 2012).

[72] Hermite versus Minkowski, preprint at arXiv:1403.1457v1 [math.NT] (6 Mar 2014).

[73] M. Dutour-Sikirić, Ph. Elbaz-Vincent, A. Kupers, J. Martinet, Voronoi complexes in higher dimensions, cohomology of GL_{N}(Z) for N ≥ 8 and the triviality of K_{8}(Z), soumis ; preprint at arXiv:1910.11598, 25 Oct. 2019.

[74] J. Martinet, Families of Equiangular Lines and Lattices, preprint at arXiv:2403.09446, 14 Mar 2024.